### v1

@@ Abstract
Code optimization is a daunting task that requires a significant level of expertise from experienced programmers. This level of expertise is not sufficient when compared to the rapid development of new hardware architectures. Towards advancing the whole code optimization process, recent approaches rely on machine learning and artificial intelligence techniques. This paper introduces a new framework to decrease the complexity of code optimization. The proposed framework builds on large language models (LLMs) and reinforcement learning (RL) and enables LLMs to receive feedback from their environment (i.e., unit tests) during the fine-tuning process. We compare our framework with existing state-of-the-art models and show that it is more efficient with respect to speed and computational usage, as a result of the decrement in training steps and its applicability to models with fewer parameters. Additionally, our framework reduces the possibility of logical and syntactical errors. Toward evaluating our approach, we run several experiments on the PIE dataset using a CodeT5 language model and RRHF, a new reinforcement learning algorithm. We adopt a variety of evaluation metrics with regards to optimization quality, and speedup. The evaluation results demonstrate that the proposed framework has similar results in comparison with existing models using shorter training times and smaller pre-trained models. In particular, we accomplish an increase of $5.6\%$ and $2.2$ over the baseline models concerning the %$\%OPT$ and $SP$ metrics.
footnotetext: 1Department of Electrical and Computer Engineering, University of Southern California 2Department of Mechanical Engineering and Aeronautics, University of Patras 3Intel Labs, Intel, USA. Correspondence to: Shukai Duan <shukaidu@usc.edu>, Nikos Kanakaris <nkanakaris@upnet.gr>.
1 Introduction
@@
Developing software (codes) from plain text descriptions or from prior implementations is a highly demanding cognitive task. However, it is significantly more challenging to optimize the code for an emerging parallel heterogeneous computing architecture. Code optimization is the task of converting a given program to a more efficient version while retaining the same input and output Bunel et al. (2016). It requires either the utilization of a higher level of optimization (e.g., -O3) during compilation or expert programmers to manually refactor their code to make it more optimized for certain hardware. Both of these tasks can be daunting with the rapid advancement in hardware, which leads to an increase in development time, bug fixing, and code optimization. Thus, there has been a noticeable shift to utilizing machine learning-based solutions for tasks related to automatic code optimization, code generation, and machine programming in general Gottschlich et al. (2018b); Shojaee et al. (2023); Gottschlich et al. (2018a). Along With recent advancements in heterogeneous devices, circuits, and computing systems, researchers have aimed to utilize machine learning and AI to initiate, design, and engineer an evolving autonomous yet compact machine programming paradigm and system that seek to provide optimized codes for distributed mobile edge computing systems having strict power/energy budgets Gottschlich et al. (2018a). One of the directions machine programming can help is the challenging task of code optimization.
@@
With the advent of transformer architectures Vaswani et al. (2017), large language models (LLMs) have emerged as the default technology to perform tasks in natural language processing (NLP). Among other applications, prominent solutions that use LLMs have shown encouraging results for software-related tasks. Undoubtedly, LLMs can perform several tasks related to programming languages, including code generation, code optimization, defect detection, and code completion, to name a few Nijkamp et al. (2023b); Wang et al. (2021). This is so easily achievable due to the large amounts of source code available in online repositories such as GitHub, which facilitate the training process Lu et al. (2021). However, even if LLMs are capable of producing code that superficially looks correct, they are unable to check its logical and syntactical validity without any external assistance. As a result, the suggested snippets of code are not always guaranteed to work as expected Husain et al. (2019). To that end, the combination of LLMs and Reinforcement Learning (RL) has been recently proposed in the literature Shojaee et al. (2023). Briefly, the inclusion of RL techniques enables LLMs to interact with their environment and receive valuable feedback. Such feedback often comes from the execution of unit tests, where the functional correctness of a given program is confirmed.
@@
Despite the radical changes and improvements that the combination of LLMs and RL offers to the analysis of programming languages, there has not been any significant progress as far as the problem of code optimization is concerned. On the one hand, the majority of the existing approaches use general-purpose datasets for fine-tuning, which in turn reduces the ability of a model to propose highly optimized versions of a given code. On the other hand, methods focusing on code optimization do not exploit RL. Thus, they are incapable of getting feedback for errors in the produced snippets of code.
@@
To mitigate the issues mentioned above, in this work, we propose PerfRL, a novel LLM-based framework that concentrates on the task of code optimization. Our approach fuses techniques from LLMs and RL to enable the utilization of external feedback, such as feedback from unit tests. By leveraging RL techniques, LLMs are able to receive information about the validity of the generated program. This helps the whole process to (i) become faster (i.e. less training is required), (ii) use a smaller model that requires less power and (iii) generate optimized code that is more likely to be free from errors. At the same time, the performance of the produced model remains the same. To the best of our knowledge, our approach is the first that specializes in the task of code generation for code optimization, while, at the same time, incorporating feedback from unit tests with respect to the correctness of the code into its learning process.
@@
Broadly speaking, there are three main challenges related to the work of this paper: (i) How can we incorporate feedback from unit tests into the training process of an LLM model? (ii) How can we make a small (compact) model perform similarly to large models with billions (or more) of parameters? (iii) How can we train such a small model such that it generates reliable code free from errors and deals with the code optimization task?
To implement and develop our approach, we use the Python programming language and PyTorch deep learning library. To evaluate our approach, we benchmark it against a list of models that deal with the specific task. Our experimental results demonstrate a significant performance improvement of our approach against the baselines. All related code and evaluation results are openly accessible on GitHub.
@@
  • We investigate the application of RL techniques combined with LLMs to improve performance on code optimization tasks. We mention that the produced LLM model specializes in the aforementioned tasks.
@@
  • We empirically test our approach using the PIE dataset, which demonstrated the superiority of our approach compared to the sate-of-the-art baselines.
2 Related Work
@@
LLMs have demonstrated promising results as far as the task of code generation is concerned. As a result, several language models (LM) for different programming languages have been proposed in the literature Chen et al. (2021). For instance, CodeT5 Wang et al. (2021) is a general language model, which is pre-trained on an extended version of the CodeSearchNet dataset Husain et al. (2019). It builds on an encoder-decoder architecture similar to T5 Raffel et al. (2020) to learn generic text representations for programming and natural languages. The authors have fine-tuned T5 to a variety of downstream tasks (e.g. code refinement and code generation) using task-specific transfer learning and multi-task learning techniques. The evaluation results have shown that CodeT5 outperformed its counterparts with respect to the CodeXGLUE benchmark Lu et al. (2021).
@@
Another family of large language models for program synthesis is CODEGEN Nijkamp et al. (2023b; a). These language models have been trained with different numbers of parameters, ranging from 350M to 16.1B. Three datasets have been used during the training period, i.e. THEPILE, BIGQUERY and BIGPYTHON. The CODEGEN models are actually autoregressive transformers with the objective of predicting the next token similar to traditional models for natural languages Vaswani et al. (2017). Along with the models, the authors have released a multi-turn programming benchmark that assists in measuring the capacity of a given model in regard to multi-turn program synthesis. The experimental results show that the multi-step program synthesis capacity of a model is positively associated with its size.
@@
More recently, the work in Madaan et al. (2023) proposed the PIE dataset. PIE is a subset of the CodeNet Puri et al. (2021) collection of code samples. It consists of trajectories of programs, where an individual programmer starts with a slower version of a program and makes changes towards improving its performance. The authors of PIE have used it to evaluate and improve the capacity of multiple variants of models from CODEGEN family. In particular, they fine-tuned these models to suggest faster versions of a given piece of code. Furthermore, they evaluated their approach by training OpenAI’s CODEX using few-shot learning. Their evaluation results exhibit a speedup improvement of up to 2.5 for more than 25% of the test programs.
A different set of modern approaches to code generation builds on the utilization of LLMs and RL. In general, the essence of incorporating RL techniques is mostly to ensure the functional correctness of the produced program Liu et al. (2023). For instance, CodeRL Le et al. (2022) combines pre-trained models with DRL for program synthesis in order to generate a program that fulfills a problem specification. It is an extension of CodeT5 that illustrates improved learning objectives and has more parameters along with better pretraining data. In CodeRL, the trained language model is used as an actor network. A critic network is also incorporated, aiming to predict the functional correctness of the generated code and provide feedback to the actor. The authors also introduce a critical sampling strategy that enables a model to regenerate programs by taking into consideration feedback from unit tests and critic scores. The evaluation results demonstrate that CodeRL can achieve state-of-the-art performance on APPS benchmark Hendrycks et al. (2021).
The work in Shen et al. (2023) introduces the ‘Rank Responses to align Test&Teacher Feedback’ (RRTF) framework. It also presents an LLM for programming languages, namely PanGu-Coder2. The main model is trained by ranking the candidate pieces of code using feedback from test cases and other heuristic preferences. Various experiments on the HumanEval, CodeEval and LeetCode benchmarks indicate that PanGu-Coder2 can reach state-of-the-art performance.
Shojaee et al. (2023) introduced PPOCoder, a task and model agnostic framework that can be used for a variety of code generation tasks. PPOCoder fuses pre-trained language models and Proximal Policy Optimization (PPO) Schulman et al. (2017), a widely used deep RL technique. PPOCoder takes into account feedback from the compiler and unit tests accompanied by syntactic-related feedback. This fosters the model to generate better code in terms of syntax and logic. The experimental results point out that PPOCode is more effective than its baselines with regard to the syntactic and functional validity of the generated codes.
Although PPO is one of the most popular policy gradient methods for RL, a list of alternative algorithms has been proposed in the literature. The purpose of these alternatives is to avoid the disadvantages of PPO (e.g. sensitivity to hyperparameters), while also reducing the overall algorithmic complexity. More specifically, Preference Ranking Optimization (PRO) Song et al. (2023) adopts the Bradley-Terry comparison from ‘Reinforcement learning from human feedback’ (RLHF) Stiennon et al. (2020); Xue et al. (2023). It also aligns the probability ranking of the responses generated by an LLM with the preference ranking by humans. The conducted experiments suggest that PRO outperforms its counterparts by effectively aligning LLMs to human preferences. Lately, RRHF, a promising algorithm has been introduced Yuan et al. (2023). In a similar fashion to RLHF, it supports the alignment of LLMs with human preferences. The authors argue that RRHF can be easily tuned and achieve a performance comparable to that of PPO in the Anthropic’s Helpful and Harmless dataset Bai et al. (2022).
3 Problem definition
Figure 1: The problem of code optimization is seen from three different perspectives, i.e. the size of the LM (Y-axis), the inference ability of the model (Z-axis), and the tuning efficiency (X-axis). Our approach (the green dot) uses a fully specialized dataset (PIE), an LM model with a medium size (CodeT5), and has a low inference ability. LMs with a big size include those of the CodeGen family and Codex.
@@ 3 Problem definition
We consider the problem of generating optimized versions of an input code. More formally, given a set of input programs $X$, the task is to generate a set of optimized programs $\hat{X}$, for each $x\in X$. The optimized program versions should take the same input and produce the same output as their original versions.
To do so, for each $x\in X$, we generate a set of candidate programs $y\in Y$, using a sampling strategy $s\in\{\text{greedy},\text{random}\}$. Then our goal is to maximize the cost function:
@@
where $y_{best}\in{Y}$ is the best candidate, the term $\text{eq}(R,y_{best})$ measures the ability of a generated sequence with a given input to match the output of the unit test, and the term $\text{perf}(y_{best})$ measures how the performance improvement of $y_{best}$ over $x$ for unit tests.
@@
Here, $\theta^{*}$ represents the optimal set of model parameters. Since recent solutions to the problem rely mostly on LLMs, the main challenge of this problem is the fact that LLMs cannot naturally interact with their environment. As a result, they can potentially produce optimized versions of a given code that look correct superficially but contain either syntactical or logical errors. Additionally, existing solutions that build on LLMs and RL do not concentrate on the task of code optimization. Thus, they cannot perform sufficiently well and may suffer from hallucination, i.e. fabricating non-compilable meaningless pieces of code. Zhang et al. (2023). A different challenge of this problem is the ever-increasing size of the required models. These models require a sufficient amount of computing resources, which in turn results in an increment of energy consumption.
@@
Considering the challenges above, we propose a different approach to train an LLM model for the code generation problem. Instead of using solely an LLM or building a general-purpose LLM model with some RL steps, we design a framework targeting the task of code generation (see Section 4.3). Figure 1 highlights the different perspectives of the code optimization problem and how our framework differs from the existing ones.
As Figure 1 illustrates, we can perceive the problem of code optimization from three different perspectives — depending on the complexity of the RL algorithm, the size of the LM, and how relevant the used dataset with the task of code optimization. For instance, in our experiments, we use a medium-sized LM (CodeT5), the RRHF RL algorithm that has a medium complexity and a dataset (PIE) that is specialized in the considered task. However, the proposed framework can facilitate different settings and combinations of the size of the selected LM model, the complexity of the RL algorithm and the relevance of the used dataset.
Some of the research questions associated with the problem of code generation are the following:
  • How can we reduce the size of the models needed to generate optimized versions of a code?
@@
  • How can we ensure that the generated codes are reliable, produce the expected results and are free from syntactical or logical errors?
4 Proposed Approach
@@
We propose PerfRL, a reinforcement learning-based LLM framework for code performance optimization. PerfRL advances the capability of LLMs to generate optimized code that improves program runtime, while also being logically and syntactically correct. To do so, it leverages techniques from LLMs and RL. It consists of three main components: (1) fine-tuning of the LLM model (Section 4.1), (2) sample generation, (Section 4.2), and (3) reinforcement learning supervision and correction (Section 4.3).
@@
Figure 2 illustrates the architecture of the proposed approach. We thoroughly describe each of the steps in this section. During the training phase, we first fine-tune an LLMs using a dataset specialized in the task of code generation. Then, we utilize the fine-tuned model to generate optimized code versions for a given input code by employing different sampling strategies, as described in Section 4.2. We then calculate a reward value and a score for each generated code and calculate a loss value based on an RL technique (Section 4.3). Then, during inference, we use the fine-tuned LLM model to generate an optimized version for each one of the given codes.
@@
Figure 2: Overview of the PerfRL framework. (a) Training. We first fine-tune an LLM model using the whole training dataset. Then, we pass the input codes into the fine-tuned LM to generate a predefined number of optimized samples for each input program. We assign a score value to each sample and calculate its reward using the reward model. We utilize the score and reward values to calculate the $L_{rank}$ and $L_{tuning}$ loss values for RL. The final loss $L$ is calculated by combining the two previous loss values and is used to retrain the model. In that way, our framework incorporates feedback from unit tests into its training process. Thus, it is more likely to generate optimized code that is free from syntactical and logical errors, mitigating hallucinations. (b) Reward model. Depending on the status of the code (e.g. can be compiled, has a run-time error or passes all the unit tests) a different reward value is given by the reward model. (c) Inference. During inference, the final LLM is utilized to generate multiple samples of candidate source codes. These generated source codes are then evaluated using the reward model. All samples that do not receive an R4 reward are filtered out. The source codes that remain after this elimination process are considered as the optimized code.
4.1 Fine-tuning of the LLM model
@@
The first step is to fine-tune an LLM model on a dataset that is specialized in the task of code optimization. For efficiency and simplicity, in our experiments (see Section 5), we use the lightweight CodeT5 model Wang et al. (2021), but our framework can operate with a variety of LLMs, regardless of their size. In the initial paper of CodeT5, the authors use either natural language (NL) or a combination of natural and programming language (PL) as input. Particularly, they pre-train their model on tasks such as Identifier-aware denoising, Identifier Tagging, Masked Identifier Prediction and Bimodal Dual Generation. Thus, we believe that this particular kind of models is more capable of understanding NL-PL inputs. As a result, we follow an NL-PL approach to feed the input into the CodeT5 model. Meanwhile, the authors in Madaan et al. (2023) argue that a few-shot sampling strategy is beneficial for producing an optimized output. Therefore, we concatenate the action’s natural language of asking the model to improve the execution performance with the input programming language and feed them into the model.
@@
where $\theta$ is the parameter of the given LLM (in our case, CodeT5), $N$ is the number of tokens, $V$ is the vocabulary set of the tokenizer, $y$ is the embedded value of the $j$-th token in the vocabulary at position $i$ in the true output sequence, and $p_{i,j}$ is the predicted probability for the $j$-th token in the vocabulary at position $i$.
4.2 Sample Generation
To generate candidate samples of code, we employ different sampling strategies during training, testing, and validation. More specifically, we use greedy sampling and random sampling.
@@ 4.2 Sample Generation
During validation and testing (i.e. inference), we generate 4 samples using greedy sampling with beam search and return the top-2 best candidates. For a given input, we generate two candidates for evaluation. Similar to Lu et al. (2021), we consider a sample as successful, when it has a better execution time compared to the input code.
4.3 Reinforcement Learning
Our RL step builds on RRHF, which is a lightweight RL framework for tuning LLMs with feedback scores. During each RL step, our objective is to maximize the probability of generating highly rewarded pieces of code by ranking and fine-tuning the LLM model. For each RL step, we sample all the data from the training set and generate the candidate outputs. Finally, for each candidate output, we compute the score, reward and loss in order to tune the model parameter $\theta$.
4.3.1 Reward
After the code generation step, we execute each sample of code with a Python interpreter. If an error is detected, an error message is displayed. If a piece of code does not have syntax errors, we test it for logical errors using the associated unit tests on a single core. The execution time $et_{o}$ for an input code $o$ is measured during the execution, assuming that there are no runtime errors within a predefined time period ($et_{o}\leq timeout$). In a similar fashion to the reward function proposed in Le et al. (2022), for each sample $y\in Y$, the reward $r(y)$ is calculated as follows:
@@ 4.3.1 Reward
As described in the paper of RRHF Yuan et al. (2023), the reward values, which are assigned to the different code samples, are irrelevant to the computation of the loss, as long as a more desirable result is associated with a higher value (see Section 4.3).
4.3.2 Score Function
For a given generated code $x$, we have a candidate sequence $y^{(c)}$, where $0<c<B$. For each $y^{(c)}$, we compute the predicted score of the sequence as the sum of the log probability of each token divided by number of tokens $t$:
@@ 4.3.3 Minimizing the Probability of Less Rewarded Outputs
4.3.4 Fine-tuning loss to maximize the best-rewarded candidate
@@
4.3.5 Loss
We sample all the input data and calculate the loss $L$ as a combination of $L_{rank}$ and $L_{tuning}$ for the samples from the same input.

### v2

@@ Abstract
Code optimization is a challenging task requiring a substantial level of expertise from developers. Nonetheless, this level of human capacity is not sufficient considering the rapid evolution of new hardware architectures and software environments. In light of this, recent research proposes adopting machine learning and artificial intelligence techniques to automate the code optimization process. In this paper, we introduce PerfRL, an innovative framework designed to tackle the problem of code optimization. Our framework leverages the capabilities of small language models (SLMs) and reinforcement learning (RL), facilitating a system where SLMs can assimilate feedback from their environment during the fine-tuning phase, notably through unit tests. When benchmarked against existing models, PerfRL demonstrates superior efficiency in terms of speed and computational resource usage, attributed to its reduced need for training steps and its compatibility with SLMs. Furthermore, it substantially diminishes the risk of logical and syntactical errors. To evaluate our framework, we conduct experiments on the PIE dataset using a lightweight large language model (i.e., CodeT5) and a new reinforcement learning algorithm, namely RRHF. For evaluation purposes, we use a list of evaluation metrics related to optimization quality and speedup. The evaluation results show that our approach achieves similar or better results compared to state-of-the-art models using shorter training times and smaller pre-trained models.
Index Terms:
code optimization, reinforcement learning, large and small language models, machine programming, programming languages, energy-efficient code optimization
I Introduction
@@
Code optimization is the task of converting a given program to a more efficient version while retaining the same input and output [1]. It requires either the utilization of a higher level of optimization (e.g., -O3) during compilation or expert programmers to manually refactor their code to make it more optimized for certain hardware. Both of these tasks can be daunting with the rapid advancement in hardware, which leads to an increase in development time, bug fixing, and code optimization. Moreover, it is significantly more challenging to optimize the code for certain environments or settings. These include emerging parallel heterogeneous computing architectures, medical devices, hard-to-access communication hubs, environments requiring dynamic software updating, new software that has to run on old or new hardware and distributed mobile edge computing systems having strict power/energy budgets [2]. As a result, there has been a noticeable shift to utilizing machine learning solutions and particularly natural language processing (NLP) techniques to automate code optimization and code generation [3, 2, 4].
@@
With the advent of transformer architectures [5, 6], language models have emerged as the default technology to perform NLP tasks. Among other applications, prominent solutions that use language models have shown encouraging results for software-related tasks, including code generation, code optimization, defect detection, and code completion, to name a few [7, 8, 9]. This is so easily achievable due to the large amounts of source code available in online repositories such as GitHub, which facilitate the training process [10]. However, even if language models are capable of producing code that superficially looks correct, they are unable to check its logical and syntactical validity without any external assistance. As a result, the suggested snippets of code are not always guaranteed to work as expected [11]. To that end, the combination of language models and Reinforcement Learning (RL) has been recently proposed in the literature [12]. Briefly, the inclusion of RL techniques enables language models to interact with their environment and receive valuable feedback. Such feedback often comes from the execution of unit tests, where the functional correctness of a given program is confirmed.
@@
Despite the radical changes and improvements that the combination of LLMs and RL offers to the analysis of programming languages, there has not been any significant progress as far as the problem of code optimization with SLMs is concerned. On the one hand, the majority of the existing approaches use general-purpose datasets for fine-tuning, which in turn reduces the ability of a model to propose highly optimized versions of a given code. On the other hand, methods focusing on code optimization do not exploit RL. Thus, they are incapable of getting feedback for errors in the produced snippets of code. Moreover, most prevailing approaches rely on LLMs whose significant demand for computation resources and high energy consumption during training and inference render them inadequate for low-cost and edge devices.
@@
To mitigate the issues mentioned above, in this work, we propose PerfRL, a novel SLM-based framework that concentrates on the task of code optimization. Our approach fuses techniques from SLMs and RL to enable the utilization of external feedback, such as feedback from unit tests. By leveraging RL techniques, SLMs are able to receive information about the validity of the generated program. This helps the whole process to (i) become faster (i.e. less training is required), (ii) use a smaller model that requires less power and fits in low-resource devices and (iii) generate optimized code that is more likely to be free from errors. At the same time, the performance of the produced model remains the same. Although RL and SLMs as algorithmic strategies have been proposed before, this is the first time that they are both combined and used for the code optimization problem.
@@
Challenges. Broadly speaking, there are four main challenges related to the work of this paper: (i) How can we incorporate feedback from unit tests into the training process of an SLM model? (ii) How can we make an SLM perform similarly to LLMs with billions (or more) of parameters? (iii) How can we train such an SLM such that it generates reliable code free from errors and deals with the code optimization task? (iv) How can we allow edge computing, cyber-physical systems and hard-to-access devices to facilitate SLMs specialized in code optimization?
Motivation. When it comes to environments that require limited resources and prioritize energy efficiency, Large Language Models (LLMs) struggle to operate effectively. These conditions are particularly relevant to cyber-physical systems, edge computing, and, more recently, the Internet of Battlefield Things (IoBT). These systems and devices undergo rapid evolution and changes, both in terms of hardware and software. As such, the need for code optimization and software integration, which supports this optimization, is of vital importance. In light of this, we argue that our proposed framework represents a groundbreaking approach to SLMs for code optimization. In fact, it democratizes AI systems research by eliminating the need to train models with an extensive number of parameters, which often leads to significant expenses.
@@
  • We propose a novel framework that trains SLMs suitable for remote devices, cyber-physical systems, edge computing or devices with limited hardware resources that require low energy consumption.
@@
  • We empirically test our approach using the PIE dataset, which demonstrated the superiority of our approach compared to the state-of-the-art baselines.
II Related Work
@@
LLMs have demonstrated promising results as far as the task of code generation is concerned. As a result, several language models (LM) for different programming languages have been proposed in the literature [13]. For instance, CodeT5 [8] is a general language model, which is pre-trained on an extended version of the CodeSearchNet dataset [11]. It builds on an encoder-decoder architecture similar to T5 [14] to learn generic text representations for programming and natural languages.
@@
Another family of large language models for program synthesis is CODEGEN [7, 15]. These language models have been trained with different numbers of parameters, ranging from 350M to 16.1B. Three datasets have been used during the training period, i.e. THEPILE, BIGQUERY and BIGPYTHON. The CODEGEN models are actually autoregressive transformers with the objective of predicting the next token similar to traditional models for natural languages [5].
@@
More recently, the work in [16] proposed the PIE dataset. PIE is a subset of the CodeNet [17] collection of code samples. It consists of trajectories of programs, where an individual programmer starts with a slower version of a program and makes changes towards improving its performance. The authors of PIE have used it to evaluate and improve the capacity of multiple variants of models from CODEGEN family. In particular, they fine-tuned these models to suggest faster versions of a given piece of code.
III Problem definition
@@
Code optimization. We consider the problem of generating optimized versions of an input code. Given a set of input programs $X$, the task is to generate a set of optimized programs $\hat{X}$ $\forall x\in X$. The optimized program versions take the same input and produce the same output as their original versions. For each $x\in X$, we generate a set of candidate programs $y\in Y$, using a sampling strategy $s\in\{\text{greedy},\text{random}\}$. Our goal is to maximize the cost function:
@@
where $y_{best}\in{Y}$ is the best candidate, the term $\text{eq}(R,y_{best})$ measures the ability of a generated sequence with a given input to match the output of the unit test and the term $\text{perf}(y_{best})$ measures how the performance improvement of $y_{best}$ over $x$ for unit tests.
@@
Here, $\theta^{*}$ represents the optimal set of model parameters. Since recent solutions to the problem rely mostly on LLMs, the main challenge of this problem is the fact that LLMs cannot naturally interact with their environment. As a result, they can potentially produce optimized versions of a given code that look correct superficially but contain either syntactical or logical errors. Additionally, existing solutions that build on LLMs and RL do not concentrate on the task of code optimization. Thus, they cannot perform sufficiently well and may suffer from hallucinations. A different challenge of this problem is the ever-increasing size of the required models. These models require a sufficient amount of computing resources, which in turn results in an increment of energy consumption.
@@
Considering the challenges above, we propose a different approach to train an LLM model for the code generation problem. Instead of using solely an LLM or building a general-purpose LLM model with some RL steps, we design a framework targeting the task of code generation. The proposed framework can facilitate different settings and combinations of the size of the selected LM model, the complexity of the RL algorithm and the relevance of the used dataset.
@@
Research questions. The research questions associated with the problem of code generation are the following: (i) How can we reduce the size of the models needed to generate optimized versions of code? (ii) How can we ensure that the generated codes are reliable, produce the expected results and are free from syntactical or logical errors?
IV Proposed Approach
@@
We propose PerfRL, a reinforcement learning-based SLM framework for code performance optimization. PerfRL advances the capability of SLMs to generate optimized code that improves program runtime, while also being logically and syntactically correct. To do so, it leverages techniques from SLMs and RL. It consists of three main components: (1) fine-tuning of the SLM model, (2) sample generation and (3) reinforcement learning supervision and correction.
@@
Figure 1 illustrates the architecture of the proposed approach. We thoroughly describe each of the steps in this section. During the training phase, we first fine-tune an SLM using a dataset specialized in the task of code generation. Then, we utilize the fine-tuned model to generate optimized code versions for a given input code by employing different sampling strategies. We then calculate a reward value and a score for each generated code and calculate a loss value based on an RL technique. Then, during inference, we use the fine-tuned SLM model to generate an optimized version for each one of the given codes.
@@
Figure 1: Overview of the PerfRL framework. (a) Training. We first fine-tune an SLM model using the whole training dataset. Then, we pass the input codes into the fine-tuned LM to generate a predefined number of optimized samples for each input program. We assign a score value to each sample and calculate its reward using the reward model. We utilize the score and reward values to calculate the $L_{rank}$ and $L_{tuning}$ loss values for RL. The final loss $L$ is calculated by combining the two previous loss values and is used to retrain the model. In that way, our framework incorporates feedback from unit tests into its training process. Thus, it is more likely to generate optimized code that is free from syntactical and logical errors, mitigating hallucinations. (b) Reward model. Depending on the status of the code (e.g. can be compiled, has a run-time error or passes all the unit tests) a different reward value is given by the reward model. (c) Inference. During inference, the final SLM is utilized to generate multiple samples of candidate source codes. These generated source codes are then evaluated using the reward model. All samples that do not receive an R4 reward are filtered out. The source codes that remain after this elimination process are considered as the optimized code.
@@
Fine-tuning of the SLM model. The first step is to fine-tune an SLM model on a dataset that is specialized in the task of code optimization. For efficiency and simplicity, in our experiments (see Section V), we use the lightweight CodeT5 model [8], but our framework can operate with a variety of SLMs and LLMs, regardless of their size. In the initial paper of CodeT5, the authors use either natural language (NL) or a combination of natural and programming language (PL) as input. Particularly, they pre-train their model on tasks such as Identifier-aware denoising, Identifier Tagging, Masked Identifier Prediction and Bimodal Dual Generation. Thus, we believe that this particular kind of models is more capable of understanding NL-PL inputs. As a result, we follow an NL-PL approach to feed the input into the CodeT5 model. Meanwhile, the authors in [16] argue that a few-shot sampling strategy is beneficial for producing an optimized output. Therefore, we concatenate the action’s natural language of asking the model to improve the execution performance with the input programming language and feed them into the model.
@@
where $\theta$ is the parameter of the given SLM (in our case, CodeT5), $N$ is the number of tokens, $V$ is the vocabulary set of the tokenizer, $y$ is the embedded value of the $j$-th token in the vocabulary at position $i$ in the true output sequence, and $p_{i,j}$ is the predicted probability for the $j$-th token in the vocabulary at position $i$.
Sample Generation. To generate candidate samples of code, we employ different sampling strategies during training, testing, and validation. More specifically, we use greedy sampling and random sampling.
@@
During validation and testing (i.e. inference), we generate 4 samples using greedy sampling with beam search and return the top-2 best candidates. For a given input, we generate two candidates for evaluation. Similar to [10], we consider a sample as successful, when it has a better execution time compared to the input code. We note that existing approaches do not incorporate into their sample strategy the aforementioned heuristic, i.e. having at least one correct sample. Thus, they are more prone to learn at least one correct source code in each step.
Reinforcement Learning. Our RL step builds on RRHF, which is a lightweight RL framework for tuning SLMs with feedback scores. During each RL step, our objective is to maximize the probability of generating highly rewarded pieces of code by ranking and fine-tuning the SLM model. For each RL step, we sample all the data from the training set and generate the candidate outputs. Finally, for each candidate output, we compute the score, reward and loss in order to tune the model parameter $\theta$.
Reward. After the code generation step, we execute each sample of code with a Python interpreter. If an error is detected, an error message is displayed. If a piece of code does not have syntax errors, we test it for logical errors using the associated unit tests on a single core. The execution time $et_{o}$ for an input code $o$ is measured during the execution, assuming that there are no runtime errors within a predefined time period ($et_{o}\leq timeout$). In a similar fashion to the reward function proposed in [18], for each sample $y\in Y$, the reward $r(y)$ is calculated as follows:
@@
As described in the paper of RRHF [19], the reward values, which are assigned to the different code samples, are irrelevant to the computation of the loss, as long as a more desirable result is associated with a higher value.
Score Function. For a given generated code $x$, we have a candidate sequence $y^{(c)}$, where $0<c<B$. For each $y^{(c)}$, we compute the predicted score of the sequence as the sum of the log probability of each token divided by number of tokens $t$:
@@
Loss. We sample all the input data and calculate the loss $L$ as a combination of $L_{rank}$ and $L_{tuning}$ for the samples from the same input.